# Convergent Privacy Framework with Contractive GNN Layers for Multi-hop Aggregations

### v1

@@ 5.3. Edge- and Node-Level Privacy of $\mathsf{CARIBOU}$
[…]
Theorem 11 ($\mathsf{CARIBOU}$’s Privacy Guarantee).
For any $\alpha>1$, let DAE training and CM training satisfy $(\alpha,\epsilon_{\mathsf{DAE}}(\alpha))$-RDP and $(\alpha,\epsilon_{\mathsf{CM}}(\alpha))$-RDP, respectively. Then, for the maximum hop $K\geq 1$, privacy failure probability $10<\delta<$ 1, Lipschitz constant $10<C_{\mathsf{L}}<$ 1, and noise variance $\sigma^{2}$, $\mathsf{CARIBOU}$ satisfies $(\epsilon_{\mathsf{CARIBOU}},\delta)$-DP, where $\epsilon_{\mathsf{CARIBOU}}=\epsilon_{\mathsf{DAE}}(\alpha)+\epsilon_{\mathsf{CM}}(\alpha)+\frac{\alpha}{2}\frac{\Delta^{2}}{\sigma^{2}}\min\left\{K,\frac{1-C_{\mathsf{L}}^{K}}{1+C_{\mathsf{L}}^{K}}\frac{1+C_{\mathsf{L}}}{1-C_{\mathsf{L}}}\right\}+\frac{\log(1/\delta)}{\alpha-1}.$ Here, $\Delta=\Delta_{e}(\mathsf{CGL})$ from Theorem 8 for edge-level privacy, or $\Delta=\Delta_{n}(\mathsf{CGL})$ from Theorem 9 for node-level privacy.
Table 3. Comparison of classification accuracy. The best accuracy and the second-best accuracy are highlighted, respectively. The symbol
[…]
Dataset Chain-S Chain-M Chain-L Chain-X Computers Facebook PubMed Cora Photo $\epsilon=1$ $\mathsf{CARIBOU}$ $84.4\%$  $82.5\%$  $70.0\%\mathrel{\hskip 5.12128pt}$ $66.0\%\mathrel{\hskip 5.12128pt}$ $92.0\%$ $74.0\%\mathrel{\hskip 5.12128pt}$ $88.2\%$ $85.1\%\mathrel{\hskip 5.12128pt}$ $95.8\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $88.0\%$ $60.6\%\mathrel{\hskip 5.12128pt}$ $88.3\%$ $75.5\%\mathrel{\hskip 5.12128pt}$ $92.5\%$ GAP $65.6\%\mathrel{\hskip 5.12128pt}$ $67.5\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $67.0\%\mathrel{\hskip 5.12128pt}$ $87.2\%$ $68.5\%\mathrel{\hskip 5.12128pt}$ $87.4\%$ $77.1\%\mathrel{\hskip 5.12128pt}$ $93.0\%$ SAGE $53.1\%$blacktriangledown $45.0\%$blacktriangledown $53.3\%$blacktriangledown $51.0\%$blacktriangledown $77.8\%$ $48.2\%$blacktriangledown $85.0\%$ $60.0\%\mathrel{\hskip 5.12128pt}$ $82.4\%$ $\epsilon=2$ $\mathsf{CARIBOU}$ $90.6\%$ $82.5\%$ $73.3\%$ $68.0\%\mathrel{\hskip 5.12128pt}$ $92.1\%$ $73.8\%\mathrel{\hskip 5.12128pt}$ $89.1\%$ $86.9\%\mathrel{\hskip 5.12128pt}$ $95.9\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $88.2\%$ $66.7\%\mathrel{\hskip 5.12128pt}$ $88.2\%$ $77.5\%\mathrel{\hskip 5.12128pt}$ $93.3\%$ GAP $65.6\%\mathrel{\hskip 5.12128pt}$ $67.5\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $68.0\%\mathrel{\hskip 5.12128pt}$ $88.3\%$ $71.7\%\mathrel{\hskip 5.12128pt}$ $87.5\%$ $77.5\%\mathrel{\hskip 5.12128pt}$ $93.4\%$ SAGE $43.8\%$blacktriangledown $45.0\%$blacktriangledown $53.3\%$blacktriangledown $51.0\%$blacktriangledown $76.1\%$ $48.1\%$blacktriangledown $84.6\%$ $60.1\%\mathrel{\hskip 5.12128pt}$ $82.4\%$ $\epsilon=4$ $\mathsf{CARIBOU}$ $90.6\%$ $82.5\%$ $73.3\%$ $69.0\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $74.0\%\mathrel{\hskip 5.12128pt}$ $89.5\%$ $87.3\%$ $95.9\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $88.9\%$ $73.3\%\mathrel{\hskip 5.12128pt}$ $88.4\%$ $75.1\%\mathrel{\hskip 5.12128pt}$ $94.2\%$ GAP $65.6\%\mathrel{\hskip 5.12128pt}$ $67.5\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $67.0\%\mathrel{\hskip 5.12128pt}$ $87.2\%$ $68.5\%\mathrel{\hskip 5.12128pt}$ $87.4\%$ $77.1\%\mathrel{\hskip 5.12128pt}$ $93.0\%$ SAGE $50.0\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $54.0\%$blacktriangledown $79.1\%$ $50.3\%$blacktriangledown $85.8\%$ $63.3\%\mathrel{\hskip 5.12128pt}$ $85.7\%$ $\epsilon=8$ $\mathsf{CARIBOU}$ $93.8\%$ $82.5\%$ $75.0\%$ $70.0\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $74.4\%\mathrel{\hskip 5.12128pt}$ $89.8\%$ $88.4\%\mathrel{\hskip 5.12128pt}$ $96.0\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $89.4\%$ $78.6\%\mathrel{\hskip 5.12128pt}$ $88.6\%$ $76.0\%\mathrel{\hskip 5.12128pt}$ $94.6\%$ GAP $65.6\%\mathrel{\hskip 5.12128pt}$ $60.0\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $67.0\%\mathrel{\hskip 5.12128pt}$ $90.1\%$ $75.0\%\mathrel{\hskip 5.12128pt}$ $88.1\%$ $79.0\%\mathrel{\hskip 5.12128pt}$ $94.6\%$ SAGE $43.8\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $61.0\%\mathrel{\hskip 5.12128pt}$ $87.9\%$ $75.6\%\mathrel{\hskip 5.12128pt}$ $84.8\%$ $75.7\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $\epsilon=16$ $\mathsf{CARIBOU}$ $93.8\%$ $85.0\%$ $73.3\%$ $73.0\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $74.5\%\mathrel{\hskip 5.12128pt}$ $89.8\%$ $88.6\%\mathrel{\hskip 5.12128pt}$ $96.0\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $40.0\%$blacktriangledown $90.4\%$ $81.2\%\mathrel{\hskip 5.12128pt}$ $88.9\%$ $77.7\%\mathrel{\hskip 5.12128pt}$ $93.8\%$ GAP $62.5\%\mathrel{\hskip 5.12128pt}$ $70.0\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $68.0\%\mathrel{\hskip 5.12128pt}$ $90.1\%$ $76.0\%\mathrel{\hskip 5.12128pt}$ $88.5\%$ $81.7\%\mathrel{\hskip 5.12128pt}$ $94.6\%$ SAGE $43.8\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $51.0\%$blacktriangledown $90.9\%$ $79.5\%\mathrel{\hskip 5.12128pt}$ $87.6\%$ $84.7\%\mathrel{\hskip 5.12128pt}$ $94.1\%$ $\epsilon=32$ $\mathsf{CARIBOU}$ $93.8\%$ $82.5\%$ $78.3\%$ $73.0\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $74.0\%\mathrel{\hskip 5.12128pt}$ $89.9\%$ $88.6\%\mathrel{\hskip 5.12128pt}$ $95.9\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $40.0\%$blacktriangledown $91.3\%$ $82.9\%\mathrel{\hskip 5.12128pt}$ $88.8\%$ $79.9\%\mathrel{\hskip 5.12128pt}$ $94.3\%$ GAP $59.4\%\mathrel{\hskip 5.12128pt}$ $65.0\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $68.0\%\mathrel{\hskip 5.12128pt}$ $90.6\%$ $77.0\%\mathrel{\hskip 5.12128pt}$ $88.8\%$ $82.5\%\mathrel{\hskip 5.12128pt}$ $94.8\%$ SAGE $43.8\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $51.0\%$blacktriangledown $90.6\%$ $79.9\%\mathrel{\hskip 5.12128pt}$ $86.9\%$ $85.2\%\mathrel{\hskip 5.12128pt}$ $94.4\%$ $\mathsf{CARIBOU}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $92.4\%$ $79.0\%\mathrel{\hskip 5.12128pt}$ $89.9\%$ $89.3\%\mathrel{\hskip 5.12128pt}$ $96.0\%$ DPDGC $59.4\%\mathrel{\hskip 5.12128pt}$ $77.5\%\mathrel{\hskip 5.12128pt}$ $63.3\%\mathrel{\hskip 5.12128pt}$ $73.0\%\mathrel{\hskip 5.12128pt}$ $92.8\%$ $86.4\%\mathrel{\hskip 5.12128pt}$ $88.1\%$ $83.9\%\mathrel{\hskip 5.12128pt}$ $96.2\%$ Plain GAP $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $91.0\%$ $79.0\%\mathrel{\hskip 5.12128pt}$ $89.3\%$ $85.2\%\mathrel{\hskip 5.12128pt}$ $95.5\%$ $\epsilon=\infty$ SAGE $59.4\%\mathrel{\hskip 5.12128pt}$ $55.0\%$blacktriangledown $60.0\%\mathrel{\hskip 5.12128pt}$ $58.0\%\mathrel{\hskip 5.12128pt}$ $91.6\%$ $79.7\%\mathrel{\hskip 5.12128pt}$ $87.0\%$ $82.1\%\mathrel{\hskip 5.12128pt}$ $94.2\%$ MLP $40.6\%$blacktriangledown $47.5\%$blacktriangledown $46.7\%$blacktriangledown $51.0\%$blacktriangledown $83.9\%$ $51.1\%$blacktriangledown $84.9\%$ $72.9\%\mathrel{\hskip 5.12128pt}$ $91.1\%$
6. Experimental Evaluation
@@
$\mathsf{CARIBOU}$ was tested over nine graph datasets. Five of the datasets have been broadly used to evaluate GNN methods, including Photo and Computers (Shchur et al., 2018), Cora and PubMed (Sen et al., 2008), Facebook (Traud et al., 2011). We also adjusted the synthetic chain-structured dataset developed under (Gu et al., 2020) into various scales, termed Chain-S, Chain-M, Chain-L and Chain-X. The chain-structured dataset has been used to understand the relations between privacy/utility and hops, as described in Section 3.1. It is considered as an important benchmark to assist the development of long-range interaction GNNs by the ML community (Dwivedi et al., 2022). More details on datasets are specialized in Appendix D.1.
Baselines.
We compare $\mathsf{CARIBOU}$ with multiple edge-level private GNNs and vanilla GNNs (Chien et al., 2023; Sajadmanesh et al., 2023; Wu et al., 2022b; Daigavane et al., 2021). In particular, we use GAP (Sajadmanesh et al., 2023) and DPDGC (Chien et al., 2023) as the baseline private GNN methods, as both have open-source implementations. SAGE and MLPare non-private baseline methods and we adapt them to the private settings. More details of the baseline methods are provided in Appendix D.2.
Model Architectures & Configurations of $\mathsf{CARIBOU}$.
We concatenate $K$ layers ranging from $1$ to $20$, with $K,\alpha_{1},\beta,C_{\mathsf{L}}$ as the hyper-parameters. If the range between two interactive nodes is long, we use larger values to test the influence of hops. Specifically, for Chain-S  and Chain-M  datasets, $\mathsf{CARIBOU}$ uses the number of hops $K\in\{15,10,9,8,7,6,5,4,3,2,1\}$, while for Chain-S  and Chain-M  datasets consisting of longer chains, $\mathsf{CARIBOU}$ takes $K\in\{20,17,15,\allowbreak 13,11,9,7,5\}$. As for standard datasets listed in Table 5, $K$ is chosen from the union sets of the aforementioned two sets, i.e., $\{20,17,15,\allowbreak 13,11,10,9,8,7,6,5,4,3,2,1\}$. To align with GAP’s configuration (Sajadmanesh et al., 2023), we set the number of hidden units to $\{16,64\}$ and use the SeLU activation function (Klambauer et al., 2017) at every layer. We adopt the Adam optimizer over $100$ epochs with a learning rate $0.001$, and pick the best accuracy to report.
Privacy Configuration and Parameters.
We implement convergent privacy allocation based on GAP’s privacy budget accounting mechanism (Sajadmanesh et al., 2023) and numerically calibrate noise level $\sigma$ by setting $\epsilon$. All edge/node-level private algorithms adopt the Gaussian mechanism to sample the i.i.d. noise to realize the desired $(\epsilon,\delta)$-DP. For comprehensive evaluation, we consider different choices of $\epsilon\in\{1,2,4,8,16,32\}$, while $\delta$ is set to be smaller than the inverse number of edges for EDP or the inverse number of node for NDP.
Hardware and software.
Our experiments have been conducted on the Ubuntu 20.04.2 LTS server, with AMD Ryzen Threadripper 3970X 32-core CPUs of 256 GB CPU memory and NVIDIA GeForce RTX 3090 of 24GB memory. $\mathsf{CARIBOU}$ equipped with edge- and node-level privacy has been implemented in PyTorch using PyTorch-Geometric (PyG) (Fey and Lenssen, 2019) framework. DP implementation adopts the autodp library (Wang, 2025), which includes analytical moments accountant and private training with DP-SGD (Abadi et al., 2016).
Table 4. Privacy Auditing via Linkteller and G-MIA. AUC score is reported.
Dataset Linkteller G-MIA ϵ=infitalic-$\epsilon=\inf$ $\epsilon=1$ $\epsilon=2$ $\epsilon=4$ $\epsilon=8$ $\epsilon=16$ $\epsilon=32$ ϵ=infitalic-$\epsilon=\inf$ $\epsilon=1$ $\epsilon=2$ $\epsilon=4$ $\epsilon=8$ $\epsilon=16$ $\epsilon=32$ Facebook $0.977$ $0.463$ $0.483$ $0.482$ $0.478$ $0.472$ $0.462$ $0.567$ $0.587$ $0.587$ $0.587$ $0.583$ $0.587$ $0.588$ PubMed $0.981$ $0.446$ $0.442$ $0.441$ $0.445$ $0.442$ $0.449$ $0.600$ $0.599$ $0.598$ $0.598$ $0.601$ $0.601$ $0.605$ Cora $0.998$ $0.427$ $0.448$ $0.399$ $0.443$ $0.451$ $0.450$ $0.645$ $0.500$ $0.500$ $0.500$ $0.500$ $0.500$ $0.500$ Photo $0.962$ $0.475$ $0.404$ $0.421$ $0.428$ $0.417$ $0.434$ $0.678$ $0.677$ $0.682$ $0.682$ $0.678$ $0.672$ $0.676$ Computers $0.860$ $0.367$ $0.364$ $0.372$ $0.363$ $0.361$ $0.384$ $0.702$ $0.701$ $0.661$ $0.707$ $0.708$ $0.711$ $0.701$
Computational overhead.
@@
We elaborate the overhead analysis, including latency and memory, in Appendix E.1.
6.1. Trade-offs of Privacy and Accuracy
@@
We first compare $\mathsf{CARIBOU}$ with the baseline methods on all datasets for their downstream classification tasks. We focus on EDP here as this is our primary privacy setting and leave the results of NDP in Appendix E.2. We run each model $3$ times for each group of hyper-parameters and report the top classification accuracy in Table 3. The mean accuracy over the 3 runs is shown in Table 9 of Appendix.
First, we highlight the results on the chain-structured datasets. In summary, we observe that the non-private versions (i.e., ϵ=infitalic-$\epsilon=\inf$) of GAP and $\mathsf{CARIBOU}$ can achieve perfect classification accuracy, while DPDGC, SAGE and MLP show relatively low accuracy. The root cause is that features of a distant node can only be learned through a sufficient number of hops of message passing, i.e., aggregating and passing features of the neighboring node one by one, but DPDGC and SAGE are not designed to leverage long-hop interactions. For MLP, as the graph structure is not utilized at all, the result is close to random guessing. After adding DP noises to protect edge-level memberships, the accuracy of both GAP and $\mathsf{CARIBOU}$ drop, but the accuracy loss of GAP is more prominent (from $100\%$ down to $57.8\%$-$62.5\%$). GAP is able to maintain reasonable accuracy even with fairly tight privacy budget $\epsilon=1$ on small chain-structured datasets like Chain-S and Chain-M. On the other hand, Chain-L and Chain-X have relatively low accuracy even for $\mathsf{CARIBOU}$ (e.g., $70.0\%$ and $66.0\%$ when $\epsilon=1$). This observed phenomenon is attributed to the increased difficulty of GNN classification with longer chains, as DP noise is injected at every layer, compounding its impact.
Regarding the other standard graph datasets, for Computers, PubMed, Cora and Photo, $\mathsf{CARIBOU}$ can outperform all the other baselines in most cases with varying $\epsilon$. It is also worthy noting that $\mathsf{CARIBOU}$ achieves fairly stable accuracy with with varying $\epsilon$ (e.g., $92.0\%$ to $92.4\%$ for Computers with $\epsilon$ growing from $1$ to $\inf$), suggesting the convergent privacy design is indeed effective. For Facebook, we found $\mathsf{CARIBOU}$ outperforms the other baselines with small $\epsilon$ (i.e., $\epsilon\leq 4$), but falls behind at larger $\epsilon$. Yet, we argue that performance at low privacy budget is more critical for privacy-sensitive datasets like Facebook, and stable accuracy across privacy budget $\epsilon$ might be more desired.
Finding 1.
@@
While GAP demonstrates advantages over existing baselines across various chain-structured datasets and parameter regimes, $\mathsf{CARIBOU}$ further advances GAP’s performance, achieving a more favorable privacy-utility trade-off.
6.2. Privacy Auditing
Following the framework outlined in Section 4.4, we implement two attack methods, Linkteller (Wu et al., 2022a) and G-MIA (Olatunji et al., 2021b), for privacy auditing. Linkteller focuses on edge-level privacy in a training-then-inference scenario. During the inference phase, Linkteller infers the private edge information held by graph holder via influence analysis and exploits adversarial queries from the feature holder. We adopt G-MIA’s attacking settings of TSTF, where models have been trained on subgraph and tested on full graph. The adversary knows the whole graph ${\bm{G}}$ and all edges contained in ${\bm{G}}$ but has no access to the subgraphs used for early training.

### v2

@@ V-B Edge- and Node-Level Privacy of $\mathsf{CARIBOU}$
VI Experimental Evaluation
@@
Datasets. $\mathsf{CARIBOU}$ was tested over nine graph datasets. Five of the datasets have been broadly used to evaluate GNN methods, including Photo and Computers [corr/abs-1811-05868], Cora and PubMed [sen2008collective], Facebook [corr/abs-1102-2166]. We also adjusted the synthetic chain-structured dataset developed under [nips/GuC0SG20] into various scales, termed Chain-S, Chain-M, Chain-L and Chain-X. The chain-structured dataset has been used to understand the relations between privacy/utility and hops, as described in Section III-A. It is considered as an important benchmark to assist the development of long-range interaction GNNs by the ML community [nips/DwivediRGPWLB22]. More details on datasets, model configurations, and privacy configurations are specialized in Appendix A-D1.
Baselines. We compare $\mathsf{CARIBOU}$ with multiple baselines [nips/chien2023differentially, uss/sajadmanesh2023gap, sp/0011L0022, arxiv/DaigavaneNode21, ccs/KolluriBHS22] about edge-level private GNNs and vanilla GNNs. To our best knowledge, GAP [uss/sajadmanesh2023gap] and DPDGC [nips/chien2023differentially] are the strongest baselines for perturbed message passing under Gaussian mechanisms in both edge/node-level DP GNNs. In addition, we consider another research direction, i.e., first perturbing the graph through randomized response and then training GNNs over the perturbed graph PertGraph [sp/0011L0022]. For a comprehensive evaluation, we adopt both research lines of works as our baselines. MLP is a baseline commonly compared with GNNs to demonstrate how well GNNs utilize graph structures, as it provides a reference counterpart to which GNNs learn the only node features without graph topology. More details of the baseline methods are provided in Appendix A-D3.
@@
Computational overhead. We elaborate the overhead analysis, including latency and memory, in Appendix A-E.
VI-A Trade-offs of Privacy and Accuracy
@@
We first compare $\mathsf{CARIBOU}$ with the baseline methods on all datasets for their downstream classification tasks and report top GNN model utility of both EDP and NDP. We run each model $3$ times for each group of hyper-parameters, reporting the top classification accuracy in Table III and the mean accuracy over the 3 runs in Table XIII of Appendix A-F. For the experiments about NDP, we set the max node degree $D_{\max}$ to $20$, following the experiment setup of GAP. We leave more reference results (e.g., different maximum node degrees) in Appendix A-F.
Regarding the standard graph datasets, for Computers, PubMed, Cora and Photo, $\mathsf{CARIBOU}$ can outperform all the other baselines in most cases with varying $\epsilon$, no matter for EDP or NDP. As established in Theorem 5, NDP requires injecting more noise compared to EDP under the same privacy budget, hence, the accuracy of NDP is often lower than EDP for standard datasets. In particular, $\mathsf{CARIBOU}$ is the only framework that can surpass MLP of most cases in NDP settings, showing effective GNN learning over structural graphs. We leave more detailed analysis on EDP and NDP in Appendix A-F and ablation study for different max node degrees in Section VI-C5.
For standard benchmark datasets with informative node features, the utility of our model approaches that of non-private methods as the privacy $\epsilon$ increases. For chain-structured graphs, the learning task primarily relies on the underlying graph topology, which is more challenging. Accordingly, model utility is more sensitive to the added noise realized by perturbed message passing. This sensitivity is due to their sparse chain structure: non-zero features are present only at the first node of each chain. Information must propagate from this source, and it can be degraded by noise accumulation during propagation. In this case if the small training set, by chance, contains an imbalanced selection of nodes (e.g., sampling nodes only near the end of a chain, far from the feature source), the task becomes significantly more difficult. This can lead to higher variance in results. To ensure a fair comparison, we use the exact same data split for all models mentioned above.
In Appendix A-G, we assess whether $\mathsf{CARIBOU}$ scales with larger real-world datasets, e.g., Reddit2, and report our results.
Takeaway 1.
@@
$\mathsf{CARIBOU}$ achieves a more favorable privacy-utility trade-off than other baselines across standard graph datasets, chain-structured datasets with various parameter settings.
VI-B Privacy Auditing
Following the PAM outlined in Section IV-C, we focus on black-box, membership-based privacy audits that match the theoretical guarantees of our DP mechanisms and those of prior perturbed message-passing methods [nips/chien2023differentially, uss/sajadmanesh2023gap]. Under this threat model, LinkTeller [sp/Wulinkteller22] and G-MIA [tpsisa/OlatunjiNK21] are canonical and complementary: LinkTeller targets edge-level membership by asking whether a specific edge exists in the training graph; G-MIA targets node-level membership by deciding whether a node and its connected edges were used during training. We adopt G-MIA’s attacking settings of TSTF, where models have been trained on subgraph and tested on full graph. The adversary knows the whole graph $\bm{G}$} and all edges contained in $\bm{G}$} but has no access to the subgraphs used for early training. Both attacks (i) are specifically designed for GNNs, (ii) operate in the transductive setting considered in our analysis, (iii) require only query access to GNN models, and (iv) are publicly available and already used to evaluate DP-GNN defenses. This makes them ideal choices as mechanism-level auditing tools in $\mathsf{CARIBOU}$.